# Enhancing Memory and Imagination Consistency in Diffusion-based World Models via Linear-Time Sequence Modeling

### v1

@@ Abstract
World models are crucial for enabling agents to simulate and plan within environments, yet existing approaches struggle with long-term dependencies and inconsistent predictions. We introduce EDELINE, a novel framework that integrates diffusion models with linear-time state space modelsto enhance memory retention and temporal consistency. EDELINE employs a recurrent embedding module based on Mamba SSMs for processing unbounded sequences, a unified architecture for joint reward and termination prediction, and dynamic loss harmonization to balance multi-task learning. Our results across multiple benchmarks demonstrate EDELINE’s superiority and robustness over prior baselines in long-horizon tasks.
Machine Learning, ICML
@@ 1 Introduction
World models (Ha & Schmidhuber, 2018), which simulate environment dynamics to enable agent planning and reasoning, represent a cornerstone of modern reinforcement learning (RL). The ability to learn compressed environment representations (Hafner et al., 2024; Schrittwieser et al., 2020) enables policy optimization through imagined trajectories and improves sample efficiency (Ye et al., 2021) compared to traditional RL methods. This capability proves particularly valuable for real-world applications, such as robotics and autonomous systems. Recent advances in world modeling demonstrate remarkable success across diverse environments, including real-world tasks (Wu et al., 2022).
@@
Existing world models can be broadly categorized into two paradigms: latent-space models and generative models. Latent-space approaches, such as (Hafner et al., 2020, 2021, 2024), leverage recurrent neural networks (RNNs) or their variants to predict future states in a compressed latent space, which enables efficient policy optimization. However, this compression often results in information loss, which compromises both generality and reconstruction quality. On the other hand, generative models, particularly diffusion-based approaches like (Alonso et al., 2024), have revolutionized world modeling by producing high-fidelity visual predictions through noise-reversal processes. Nevertheless, prior generative models rely on fixed-length observation-action windows, which truncate historical context and fail to capture extended temporal dependencies. This limitation becomes particularly problematic in partially observable environments, where agents must retain and reason over long sequences of observations to make informed decisions. Moreover, the architectural separation of reward prediction, termination signals, and observation modeling in existing frameworks can lead to suboptimal representation sharing and optimization conflicts, further hindering performance.
@@
In order to mitigate the long sequence dependency issues, recent emergence of state space models (SSMs), such as (Gu et al., 2022a, b; Smith et al., 2023; Gu & Dao, 2024b, a) offer a complementary advantage as they are tailored to model long-term dependencies efficiently. With linear-time complexity and selective state updates (Gu & Dao, 2024b), SSMs can theoretically process unbounded sequences while preserving critical historical information. This capability is particularly valuable for world modeling, where accurate trajectory prediction often necessitates retaining and reasoning over extended observation-action histories.
@@
Motivated by these considerations, we introduce EDELINE (Enhancing Diffusion-basEd world models via LINEar-time sequence modeling), a unified framework that synthesizes the advantages of diffusion models and SSMs. EDELINE advances the state-of-the-art through three key innovations: (1) Memory Enhancement: A recurrent embedding module (REM) based on Mamba SSMs for processing unbounded observation-action sequences to enable adaptive memory retention beyond fixed-context limitations, (2) Unified Framework: Direct conditioning of reward and termination prediction on REM hidden states for eliminating separate networks for efficient representation sharing, and (3) Dynamic Loss Harmonization: Adaptive weighting of observation, reward, and termination losses for addressing scale disparities in multi-task optimization. To validate our approach, we provide theoretical and experimental justifications to support that hidden information can preserve enhanced detail for more accurate imaginary compared with stacked frames inputs. To further demonstrate EDELINE’s capabilities, we conduct extensive validation across various benchmarks. The experimental results reveal EDELINE’s superior performance with 1.866 human normalized scores in Atari100K, which surpasses all current model-based approaches. To substantiate EDELINE’s ability to preserve long temporal information for consistent predictions, we evaluate performance on MiniGrid and VizDoom environments. Both qualitative and quantitative results demonstrate superior temporal consistency in modeling and imaginary quality across 2D and 3D environments. Our ablation studies validate the effectiveness of each architectural component for world modeling performance. Our contributions are summarized as follows:
  1. 1.
@@
    We introduce a unified architecture that integrates a Next-Frame Predictor for future state imaginary, a Recurrent Embedding Module for temporal sequence processing, as well as a Reward/Termination Predictor.
  2. 2.
    Theoretical guarantees demonstrate that EDELINE retains strictly more information than fixed-context approaches, with advantages growing linearly over time.
  3. 3.
@@
    Experimental results in various benchmarks validate EDELINE’s precise prediction of reward-critical elements that prior models exhibit structural inaccuracies.
@@ 2.1 Diffusion Models
Diffusion models have revolutionized high-resolution image generation through their noise-reversal process. Foundational works including DDPM (Ho et al., 2020) and DDIM (Song et al., 2021a) established core principles for subsequent developments. Score-based models (Song & Ermon, 2019; Song et al., 2021b) enhanced sampling efficiency through gradient estimation of data distributions, while energy-based models (Du et al., 2023) introduced robust optimization properties through probabilistic state modeling. The application of diffusion models in reinforcement learning has expanded significantly. These models serve as policy networks for efficient offline learning (Zhendong et al., 2023; Anurag et al., 2023; Tim et al., 2023), enable diverse strategy generation in planning tasks (Michael et al., 2023; Zhixuan et al., 2023), and provide novel approaches to reward modeling (Felipe Pinto Coelho et al., 2023). MetaDiffuser (Ni et al., 2023) demonstrated their effectiveness as conditional planners in offline meta-RL, while other works have utilized diffusion models for trajectory modeling and synthetic experience generation (Cong et al., 2023).
2.2 Model-based RL and Generative Game Engines
@@
World models serve as a fundamental component in model-based RL and enable sample-efficient and safe learning through simulated environments. SimPLe (Łukasz Kaiser et al., 2020) established the groundwork by introducing world models to the Atari domain and proposing the Atari 100k benchmark. Dreamer (Hafner et al., 2020) advanced this field through reinforcement learning from latent space predictions, which DreamerV2 (Hafner et al., 2021) further refined with discrete latents to mitigate compounding errors. DreamerV3 (Hafner et al., 2024) achieved a significant milestone by demonstrating human-level performance across multiple domains with fixed hyperparameters. Recent architectural innovations have expanded world model capabilities. TWM (Robine et al., 2023) and STORM (Zhang et al., 2023) incorporated Transformer architectures for enhanced sequence modeling, while IRIS (Micheli et al., 2023) developed a discrete image token language for structured learning. R2I (Samsami et al., 2024) addressed fundamental challenges in long-term memory and credit assignment. The integration of generative modeling approaches has further advanced world model capabilities. DIAMOND (Alonso et al., 2024) marked a significant breakthrough by incorporating diffusion models, which achieved superior visual fidelity and state-of-the-art performance on the Atari 100k benchmark. This success inspired broader applications of generative approaches in interactive environments. Another line of research explores training generative world engines using pre-collected datasets instead of RL-in-the-loop learning. GameGAN (Kim et al., 2020) pioneered this direction through GAN-based environment modeling, while Genie (Bruce et al., 2024) advanced these capabilities by generating complex platformer environments from image prompts. GameNGen (Valevski et al., 2024) established new standards for visual fidelity and scalability through diffusion-based environment simulation. It is important to note that these generative world engines do not involve RL-in-the-loop training and are therefore orthogonal to our work.
2.3 State Space Models in Reinforcement Learning
@@
SSMs have demonstrated remarkable effectiveness in modeling long-term dependencies and structured dynamics. Foundational works (Gu et al., 2020, 2022a; Gupta et al., 2022; Gu et al., 2022b; Smith et al., 2023; Ramin et al., 2023; Gu & Dao, 2024b, a) introduced efficient sequence modeling approaches. Recent applications in RL include structured state space models for in-context learning (Chris et al., 2023), DecisionMamba (Huang et al., 2024)’s adaptation of Decision Transformer (Chen et al., 2021) architecture, and Drama (Wenlong et al., 2024)’s integration of SSMs in world model learning. Our work advances the state-of-the-art by synergistically combining diffusion-based world models with state space models. While previous works have explored these approaches separately, EDELINE demonstrates that their integration enables superior temporal consistency and extended imagination horizons in world model learning. This architectural innovation enhances the visual fidelity of diffusion models through the integration of SSMs’ efficient temporal processing capabilities, which addresses the key limitations of existing world model approaches.
@@ 3 Background
We provide additional background material on score-based diffusion models and multi-task world model learning in Appendix M. In this Section, we focus on the essential concepts necessary for understanding our EDELINE framework.
@@ 3.1 Reinforcement Learning and World Models
The problem considered in this study focuses on image-based reinforcement learning (RL), formulated as a Partially Observable Markov Decision Process (POMDP) (Åström, 1965) defined by tuple $(S,A,O,T,R,O,\gamma)$. Our formulation specifically considers high-dimensional image observations as inputs, as described in Section 1. The state space $S$ comprises states $s_{t}\in S$, while the action space $A$ can be either discrete or continuous with actions $a_{t}\in A$. The observation space $O$ contains image observations $o_{t}\in\mathbb{R}^{3\times H\times W}$. A transition function $T:S\times A\times S\rightarrow[0,1]$ characterizes the environment dynamics $p(s_{t+1}|s_{t},a_{t})$, while the reward function $R:S\times A\times S\rightarrow\mathbb{R}$ maps transitions to scalar rewards $r_{t}\in\mathbb{R}$. The observation function $O:S\times O\rightarrow[0,1]$ establishes observation probabilities $p(o_{t}|s_{t})$. The objective centers on learning a policy $\pi$ that maximizes the expected discounted return $\mathbb{E}_{\pi}[\sum_{t\geq 0}\gamma^{t}r_{t}]$, with discount factor $\gamma\in[0,1]$. Model-based Reinforcement Learning (MBRL) (Sutton, 1988) achieves this objective by learning a world model that encapsulates the environment dynamics $p(o_{t+1},r_{t}|o_{\leq t},a_{\leq t})$. MBRL enables learning in imagination through three systematic stages: (1) collecting real environment interactions, (2) updating the world model, and (3) training the policy through world model interactions.
@@ 3.2 Linear-Time Sequence Modeling with Mamba
SSMs (Gu et al., 2022a) provide an alternative paradigm to attention-based architectures for sequence modeling. The Mamba architecture (Gu & Dao, 2024b) introduces a selective state space model that offers linear time complexity and efficient parallel processing, which employs variable-dependent projection matrices to implement its selective mechanism, thus overcoming the inherent limitations of computational inefficiency and quadratic complexity in conventional SSMs (Gu et al., 2020, 2022a; Smith et al., 2023; Gu & Dao, 2024b). The foundational mechanism of Mamba is characterized by a linear continuous-time state space formulation via first-order differential equations as follows:
@@
where $x(t)$ represents the latent state, $u(t)$ denotes the input, and $y(t)$ indicates the output. The matrix $A$ adheres to specifications from (Gu et al., 2022b). The primary innovation compared to traditional SSMs lies in $B(u(t))$ and $C(u(t))$, which function as state-dependent linear operators to enable selective state updates based on input content. For discretization, the system employs the zero-order-hold (ZOH) rule (Chifu et al., 2018) to transform the $A$ and $B$ matrices into $\tilde{A}=\exp(\Delta A)$ and $\tilde{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B$, where the step size $\Delta$ serves as a variable-dependent parameter. This transformation enables SSMs to process continuous inputs as discrete signals and converts the original Linear Time-Invariant (LTI) equation into a recurrence format.
@@ 3.3 Diffusion-based World Model Learning
To adapt diffusion models for world modeling, which offers superior sample quality and tractable likelihood estimation, a key requirement is modeling the conditional distribution $p(o_{t+1}|o_{\leq t},a_{\leq t})$, where $o_{t}$ and $a_{t}$ represent observations and actions at time step $t$. The denoising process incorporates both the noised next observation and the conditioning context as input: $D_{\theta}(o_{t+1}^{\tau},\tau,o_{\leq t},a_{\leq t})$. While diffusion-based world models (Alonso et al., 2024) have shown promise, the state-of-the-art approach DIAMOND (Alonso et al., 2024) exhibits limitations although it achieves superior performance on the Atari 100k benchmark (Łukasz Kaiser et al., 2020). These models face two critical limitations. The first limitation stems from their constrained conditioning context, which typically considers only the most recent observations and actions. For instance, DIAMOND restricts its context to the last four observations and actions in the sequence. This constraint impairs the model’s capacity to capture long-term dependencies and leads to inaccurate predictions in scenarios that require extensive historical context. The second limitation in current diffusion-based world models lies in their architectural separation of predictive tasks. For example, DIAMOND implements a separate neural network with convolutional and recurrent layers for reward and termination prediction. This separation prevents the sharing of learned representations between the diffusion model and these predictive tasks and results in reduced overall learning efficiency of the system.

### v2

# EDELINE: Enhancing Memory in Diffusion-based World Models via Linear-Time Sequence Modeling

@@ Abstract
World models represent a promising approach for training reinforcement learning agents with significantly improved sample efficiency. While most world model methods primarily rely on sequences of discrete latent variables to model environment dynamics, this compression often neglects critical visual details essential for reinforcement learning. Recent diffusion-based world models condition generation on a fixed context length of frames to predict the next observation, using separate recurrent neural networks to model rewards and termination signals. Although this architecture effectively enhances visual fidelity, the fixed context length approach inherently limits memory capacity. In this paper, we introduce EDELINE, a unified world model architecture that integrates state space models with diffusion models. Our approach outperforms existing baselines across visually challenging Atari 100k tasks, memory-demanding Crafter benchmark, and 3D first-person ViZDoom environments, demonstrating superior performance in all these diverse challenges.
@@ 1 Introduction
World models NEURIPS2018_2de5d166 constitute a foundational element of modern reinforcement learning (RL) by simulating environment dynamics for agent planning and reasoning. The capacity to learn environment representations hafner2024DreamerV3 ; Schrittwieser_2020 facilitates policy optimization through imagined trajectories, which substantially enhances sample efficiency NEURIPS2021_d5eca8dc relative to conventional RL approaches. This capability is especially valuable for real-world applications in robotics and autonomous systems. Recent advances have shown exceptional performance across diverse environments, including real-world tasks wu2022daydreamer .
@@
Existing world models fall into two principal paradigms: latent-space models and generative models. Latent-space approaches Hafner2020Dreamer ; hafner2021DreamerV2 ; hafner2024DreamerV3 employ recurrent neural networks (RNNs) or variants to predict future states within a compressed latent space for efficient policy optimization. This compression, however, introduces information loss that compromises generality and reconstruction quality. Generative models, particularly diffusion-based approaches alonso2024diamond , have transformed world modeling through high-fidelity visual predictions via noise-reversal processes. Nevertheless, prior generative models depend on fixed-length observation-action windows that truncate historical context and fail to capture extended temporal dependencies. This limitation presents a challenge especially in partially observable environments where agents must retain and reason over prolonged observation sequences for informed decisions. Moreover, the architectural segregation of reward prediction, termination signals, and observation modeling in existing frameworks can potentially lead to suboptimal representation sharing and optimization conflicts that further impair performance.
@@
In order to mitigate long sequence dependency issues, recent state space models (SSMs) gu2022efficiently ; gu2022parameterizationinitializationdiagonalstate ; smith2023simplified ; gu2024mamba ; dao2024transformersssmsgeneralizedmodels provide a complementary advantage through their capacity to model long-term dependencies efficiently. With linear-time complexity and selective state updates gu2024mamba , SSMs can process theoretically unbounded sequences while preserving critical historical information. This capability is particularly valuable for world modeling, where accurate trajectory prediction often necessitates retention and reasoning across extended observation-action histories. In the world modeling domain, recent work such as R2I samsami2024r2i has addressed fundamental challenges in long-term memory and credit assignment, demonstrating the importance of enhanced temporal reasoning capabilities.
@@
Based on these considerations, we introduce EDELINE (Enhancing Diffusion-basEd World Models via LINEar-Time Sequence Modeling), a unified framework that integrates the advantages of diffusion models and SSMs. EDELINE advances the state-of-the-art (SOTA) through three key innovations: (1) Memory Enhancement: A recurrent embedding module (REM) based on Mamba SSMs that processes unbounded observation-action sequences to enable adaptive memory retention beyond fixed-context limitations, (2) Unified Framework: Direct conditioning of reward and termination prediction on REM hidden states that eliminates separate networks for efficient representation sharing, and (3) Dynamic Loss Harmonization: Adaptive weighting of observation and reward losses that addresses scale disparities in multi-task optimization. To validate EDELINE’s effectiveness, we conduct extensive evaluations. It achieves 1.87$\times$ human normalized scores on the sample-efficiency challenging Atari 100k Kaiser2020SimPLe benchmark and surpasses all model-based methods that do not use look-ahead search. The ablation studies confirm the effectiveness of each architectural component for world modeling performance. To substantiate EDELINE’s capacity to preserve long temporal information for consistent predictions, we evaluate performance on environments that require long-term memory capability: MiniGrid-Memory chevalier-boisvert2023minigrid , Crafter hafner2022benchmarking , and VizDoom kempka2016vizdoomdoombasedairesearch . Both qualitative and quantitative results show superior temporal consistency in modeling and imaginary quality across 2D and 3D environments. Furthermore, EDELINE shows significant performance improvements compared to prior diffusion-based world models. Our contributions can be summarized as follows:
@@
  • We introduce a unified architecture EDELINE that integrates a Next-Frame Predictor for future observation imaginary, a Recurrent Embedding Module for temporal sequence processing, and a Reward/Termination Predictor to address the long-term memory limitations in existing models.
  • EDELINE utilizes an SSMs-based embedding module that overcomes the fixed context limitations of prior diffusion-based methods and enhances performance in memory-demanding environments.
@@
  • Our experimental insights across various benchmarks validate EDELINE’s precise prediction of reward-critical elements where prior diffusion-based world models exhibit structural inaccuracies.
@@ 2.1 Diffusion Models
Diffusion models have revolutionized high-resolution image generation through their noise-reversal process. Foundational works including DDPM ho2020ddpm and DDIM song2021ddim established core principles for subsequent developments. Score-based models song2019generative ; song2021scorebased enhanced sampling efficiency through gradient estimation of data distributions, while energy-based models du2024reducereuserecyclecompositional introduced robust optimization properties via probabilistic state modeling. The application of diffusion models in RL has expanded significantly. These models serve as policy networks for efficient offline learning wang2023diffusion ; ajay2023is ; pearce2023imitating , enable diverse strategy generation in planning tasks janner2022planningdiffusionflexiblebehavior ; liang2023adaptdiffuser , and provide novel approaches to reward modeling nuti2023extracting . MetaDiffuser ni2023metadiffuserdiffusionmodelconditional showed effectiveness as conditional planners in offline meta-RL, while others have adopted diffusion models for trajectory modeling and synthetic experience generation lu2023synthetic .
2.2 World Models for Sample-Efficient RL
World models serve as a fundamental component in model-based RL, enabling sample-efficient and safe learning through simulated environments. SimPLe Kaiser2020SimPLe established the groundwork by introducing world models to the Atari domain and proposing the Atari 100k benchmark. Dreamer Hafner2020Dreamer advanced this field through RL from latent space predictions, which DreamerV2 hafner2021DreamerV2 further refined with discrete latents to mitigate compounding errors. DreamerV3 hafner2024DreamerV3 achieved a significant milestone by demonstrating human-level performance across multiple domains with fixed hyperparameters.
@@ 2.2 World Models for Sample-Efficient RL
Recent architectural innovations have expanded world model capabilities. TWM robine2023TWM and STORM zhang2023storm incorporated Transformer architectures NIPS2017_3f5ee243 for enhanced sequence modeling, while IRIS micheli2023iris developed a discrete image token language for structured learning. R2I samsami2024r2i addressed fundamental challenges in long-term memory and credit assignment. The integration of generative modeling approaches has further advanced world model capabilities. DIAMOND alonso2024diamond marked a significant breakthrough by incorporating diffusion models, which achieved superior visual fidelity and state-of-the-art (SOTA) performance on the Atari 100k benchmark. This success inspired broader applications of generative approaches in interactive environments.
2.3 Generative Game Engines
Another line of research explores training generative world engines using pre-collected datasets instead of RL-in-the-loop learning. GameGAN kim2020GameGAN pioneered this direction through GAN-based environment modeling, while Genie bruce2024genie advanced these capabilities by generating complex platformer environments from image prompts. GameNGen valevski2024diffusionmodelsrealtimegame established new standards for visual fidelity and scalability through diffusion-based environment simulation. Recently, GAIA hu2023gaia1generativeworldmodel ; russell2025gaia2controllablemultiviewgenerative and Pandora xiang2024pandorageneralworldmodel have further expanded this field by developing generative world models that leverage video, text, and action inputs to produce realistic scenarios. It is important to note that these generative world methodologies do not involve RL-in-the-loop training and are therefore orthogonal to our work.
2.4 State Space Models (SSMs) in Reinforcement Learning
@@
SSMs have demonstrated remarkable effectiveness in modeling long-term dependencies and structured dynamics. Foundational works hippo ; gu2022efficiently ; gupta2022diagonal ; gu2022parameterizationinitializationdiagonalstate ; smith2023simplified ; hasani2023liquid ; gu2024mamba ; dao2024transformersssmsgeneralizedmodels introduced efficient sequence modeling approaches. Recent applications in RL include structured SSMs for in-context learning lu2023structured , DecisionMamba huang2024decisionmamba ’s adaptation of Decision Transformer chen2021decision architecture, and Drama anonymous2025drama ’s integration of SSMs in world model learning. Our work advances the state-of-the-art by synergistically combining diffusion-based world models with SSMs. While previous works have explored these approaches separately, EDELINE demonstrates that their integration enables superior temporal consistency and extended imagination horizons in world model learning. This architectural innovation enhances the visual fidelity of diffusion models through the integration of SSMs’ efficient temporal processing capabilities, which addresses the key limitations of existing world model approaches.
@@ 3 Background
In this Section, we focus on the essential concepts necessary for understanding our EDELINE framework. We provide additional background material on score-based diffusion models and multi-task world model learning in Appendix C.
@@ 3.1 Reinforcement Learning and World Models
The problem considered in this study focuses on image-based reinforcement learning (RL), formulated as a Partially Observable Markov Decision Process (POMDP) strm1965OptimalCO defined by tuple $(S,A,O,P,R,O,\gamma)$. Our formulation specifically considers high-dimensional image observations as inputs, as described in Section 1. The state space $S$ comprises states $s_{t}\in S$, while the action space $A$ can be either discrete or continuous with actions $a_{t}\in A$. The observation space $O$ contains image observations $o_{t}\in\mathbb{R}^{3\times H\times W}$. A transition function $P:S\times A\times S\rightarrow[0,1]$ characterizes the environment dynamics $p(s_{t+1}|s_{t},a_{t})$, while the reward function $R:S\times A\times S\rightarrow\mathbb{R}$ maps transitions to scalar rewards $r_{t}\in\mathbb{R}$. The observation function $O:S\times O\rightarrow[0,1]$ establishes observation probabilities $p(o_{t}|s_{t})$. The objective centers on learning a policy $\pi$ that maximizes the expected discounted return $\mathbb{E}_{\pi}[\sum_{t\geq 0}\gamma^{t}r_{t}]$, with discount factor $\gamma\in[0,1]$. Model-based Reinforcement Learning (MBRL) Sutton1988LearningTP achieves this objective by learning a world model that encapsulates the environment dynamics $p(o_{t+1},r_{t}|o_{\leq t},a_{\leq t})$. MBRL enables learning in imagination through three systematic stages: (1) collecting real environment interactions, (2) updating the world model, and (3) training the policy through world model interactions.
@@ 3.2 Linear-Time Sequence Modeling with Mamba
SSMs gu2022efficiently provide an alternative paradigm to attention-based architectures for sequence modeling. The Mamba architecture gu2024mamba introduces a selective state space model that offers linear time complexity and efficient parallel processing, which employs variable-dependent projection matrices to implement its selective mechanism, thus overcoming the inherent limitations of computational inefficiency and quadratic complexity in conventional SSMs hippo ; gu2022efficiently ; smith2023simplified ; gu2024mamba . The foundational mechanism of Mamba is characterized by a linear continuous-time state space formulation via first-order differential equations as follows:
@@
where $x(t)$ represents the latent state, $u(t)$ denotes the input, and $y(t)$ indicates the output. The matrix $A$ adheres to specifications from gu2022parameterizationinitializationdiagonalstate . The primary innovation compared to traditional SSMs lies in $B(u(t))$ and $C(u(t))$, which function as state-dependent linear operators to enable selective state updates based on input content. For discretization, the system employs the zero-order-hold (ZOH) rule yang2018improvingclosedlooptrackingperformance to transform the $A$ and $B$ matrices into $\tilde{A}=\exp(\Delta A)$ and $\tilde{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B$, where the step size $\Delta$ serves as a variable-dependent parameter. This transformation enables SSMs to process continuous inputs as discrete signals and converts the original Linear Time-Invariant (LTI) equation into a recurrence format.
@@ 3.3 Diffusion-based World Model Learning
To adapt diffusion models for world modeling, which offers superior sample quality and tractable likelihood estimation, a key requirement is modeling the conditional distribution $p(o_{t+1}|o_{\leq t},a_{\leq t})$, where $o_{t}$ and $a_{t}$ represent observations and actions at time step $t$. The denoising process incorporates both the noised next observation and the conditioning context as input: $D_{\theta}(o_{t+1}^{\tau},\tau,o_{\leq t},a_{\leq t})$. While diffusion-based world models alonso2024diamond have shown promise, the state-of-the-art approach DIAMOND alonso2024diamond exhibits limitations although it achieves superior performance on the Atari 100k benchmark Kaiser2020SimPLe . These models face two critical limitations. The first limitation stems from their constrained conditioning context, which typically considers only the most recent observations and actions. For instance, DIAMOND restricts its context to the last four observations and actions in the sequence. This constraint impairs the model’s capacity to capture long-term dependencies and leads to inaccurate predictions in scenarios that require extensive historical context. The second limitation in current diffusion-based world models lies in their architectural separation of predictive tasks. For example, DIAMOND implements a separate recurrent neural network for reward and termination prediction. This separation prevents the sharing of learned representations between the diffusion model and these predictive tasks and results in reduced overall learning efficiency of the system.